the ratio of rates of diffusion of co2 and so2 at same temperature and pressure

Grade:12th pass

3 Answers

Khimraj
5 years ago
Hope it clears...................\sqrt{64/44} = \sqrt{16/11} result for graham`s law of diffusion So for CO2; M = 44 So for SO2; M = 64 so R1/R2 = \sqrt{M2/M1}As according to GRAHAM’s LAW; R1/R2 =  
Khimraj
5 years ago
\sqrt{M2/M1}As according to GRAHAM’s LAW; R1/R2 =  
for CO2; M1 = 44 So for SO2; M2 = 64
 R1/R2 = \sqrt{64/44} = \sqrt{16/11}
Hope it clears.........................................
 
Khimraj
5 years ago
As according to GRAHAM’s LAW; R1/R2 =  \sqrt{M2/M1}
for CO2; M1 = 44 So for SO2; M2 = 64
 R1/R2 = \sqrt{64/44} = \sqrt{16/11}
Hope it clears.........................................